The velocity of a particle is given by v =
v0
1 + At2
, where v0 = 25 m s−1
and A = 2.0 s−2
. What
is the acceleration at t = 0? At t = 2.0 s? At t → ∞?
1
Expert's answer
2019-09-16T09:51:16-0400

The velocity of a particle


"v(t)=v_0(1+At^2)=25(1+2.0t^2)"

The acceleration of a particle


"a(t)=\frac{dv(t)}{dt}=(25(1+2.0t^2))'=100t"

So


"a(0)=0\:\rm m/s^2"

"a(2.0)=200 \:\rm m/s^2"

"a(\infty)=\infty"